Reduce the homogeneous equation to the separated one

(x+2y)dx-xdy=0


1
Expert's answer
2022-02-02T09:26:16-0500

"(x+2y)dx-xdy=0\\
(x+2y)dx=xdy\\
\frac{dy}{dx}=\frac{x+2y}{x}...(1)\\"

Let "y=vx"

"\frac{dy}{dx}=v+x\frac{dv}{dx}"

From (1),

"v+x\frac{dv}{dx}=\frac{x+2vx}{x}=1+2v\\
x\frac{dv}{dx}=1+v\\
\frac{dv}{1+v}=\frac{dx}{x}\\"

Integrating both sides, we get

"ln(1+v)= lnx+ ln c\\
ln(1+v)-lnx=lnc\\
ln(\frac{1+v}{x})=lnc\\
\frac{1+v}{x}=c\\
\frac{1+\frac{y}{x}}{x}=c\\
x+y=x^2c"